OpenStudy (anonymous):

What is the equation of the line of symmetry for the graph of y = -2(x - 4)2 + 3? x = -2 x = 2 x = 3 x = 4

OpenStudy (anonymous):

is your equation y = -2(x - 4)2 + 3 or y = -2(x - 4)^2 + 3

OpenStudy (anonymous):

^2

OpenStudy (anonymous):

ok well first start by distributing

OpenStudy (anonymous):

CouldI put the x answers into the equation?

OpenStudy (anonymous):

not yet

OpenStudy (anonymous):

so what do you get after you distribute?

OpenStudy (anonymous):

y = -2x^2+8^2+3 correct?

OpenStudy (anonymous):

@badATalgebra

OpenStudy (precal):

you already have this in vertex form, your axis of symmetry is the first coordinate of your vertex